OpenStudy (anonymous):

The volume of a prism is found by multiplying the area of the base by the height. A rectangular prism has a volume of 144 cm3 , and a square base with a width of 3 cm. What is the height of the prism, in cm? A. 4 B. 8 C. 48 D. 16

OpenStudy (johnweldon1993):

Okay so we have the equation \[V = \text {area of base} \times height\] Well we see that the base...is a square with width of 3....how do we find the area of a square?

OpenStudy (johnweldon1993):

Well the area of a square is length * width....we know a square has 4 equal sides so this is simply \[\text {area of a square} = \text {length} \times \text{width} \ = 3 \times 3 = 9\] So now that we know that area AND the volume...we can solve for the height \[V = \text{area of base} \times height\] \[144 = 9 \times height\] What does height equal?

OpenStudy (anonymous):

height = 16